Tan tita+1/tan tita is 5 then find the value of tan sqare tita+1/tan square tita

Grade:8

1 Answers

poojitha
6 years ago
Tan tita+1/Tan tita=5,  Tan^2tita+1/Tan^2tita=(Tan tita+1/Tan tita)^2-(2.Tan tita.1/Tan tita)=(5)^2-2=25-2=23